Hayhayz (hayhayz):

What is the discontinuity of the function f(x) = the quantity of x squared minus 3 x minus 28, all over x plus 4? @zakkdanger

OpenStudy (misty1212):

Hi!!

Hayhayz (hayhayz):

http://prntscr.com/99lens

OpenStudy (misty1212):

hard to read, but whatever it is, set the denominator equal to zero and solve

OpenStudy (misty1212):

\[x+4=0\\ x=-4\]

OpenStudy (zakkdanger):

-4

Hayhayz (hayhayz):

thats really all you do?

OpenStudy (misty1212):

really easy right?q

Hayhayz (hayhayz):

so it would be (-4,-11)

OpenStudy (misty1212):

guess so, it is the only one with a -4 first right?

Hayhayz (hayhayz):

yeah

OpenStudy (misty1212):

i can show you how to get the -11 if you like

Hayhayz (hayhayz):

yes please

OpenStudy (misty1212):

factor and cancel

OpenStudy (zakkdanger):

Sh!t she beat me to it. GOD DAMMIT

Hayhayz (hayhayz):

lol zakk

OpenStudy (misty1212):

\[\frac{x^2-3x-28}{x+4}=\frac{(x+4)(x-7)}{x+4}=x-7\]

OpenStudy (misty1212):

then if \(x=-4\) plug it in to the cancelled answer and get \[-4-7=-11\]

OpenStudy (misty1212):

lol @jakelingo

Hayhayz (hayhayz):

Thank you! can i tag you in another one?

OpenStudy (misty1212):

sure i got ten minutes, but i bet zak can do it too

Hayhayz (hayhayz):

alright

OpenStudy (zakkdanger):

0_0 Im here

OpenStudy (zakkdanger):

What the ?

OpenStudy (misty1212):

what happened to the ladies in your picture?

OpenStudy (zakkdanger):

IDK i was prankin peeps in chat then I posted one of me dressed like Sir Charles P. Scene